OpenStudy (anonymous):

Sal currently has an account balance of $2,835.48. He opened the account five years ago with a deposit of $2,310.72. If the interest compounds twice a year, what is the interest rate on the account?

OpenStudy (dumbcow):

2835.48= 2310.72(1+r/2)^10 1.2271 = (1+r/2)^10 1.02067 = 1+ r/2 .02067 = r/2 r = 0.04135